(a) The differential equation (2π‘₯ 2 + 𝑏𝑦 2 )𝑑π‘₯ + 𝑐π‘₯𝑦 𝑑𝑦 = 0 can made exact by multiplying with integrating factor 1 π‘₯ ⁄ 2.

Then find the relation between 𝑏 and 𝑐. (b) Find one fourth roots of unity


1
Expert's answer
2022-02-14T16:48:34-0500

(a)


"(\dfrac{1}{x^2})(2x^2+by^2)dx+(\dfrac{1}{x^2})cxydy=0"

"(2+b\dfrac{y^2}{x^2})dx+c(\dfrac{y}{x})dy=0"

"\dfrac{\partial}{dx}(c(\dfrac{y}{x}))=\dfrac{\partial}{\partial y}(2+b\dfrac{y^2}{x^2})"

"-c(\dfrac{y}{x^2})=2b(\dfrac{y}{x^2})"

"c=-2b"

(b) The polar form ofΒ "1" Β isΒ "\cos(0)+i\sin(0)."


"k=0, \sqrt[4]{1}(\cos(\dfrac{0+2\pi (0)}{4})+i\sin(\dfrac{0+2\pi (0)}{4}))=1"

"k=1, \sqrt[4]{1}(\cos(\dfrac{0+2\pi (1)}{4})+i\sin(\dfrac{0+2\pi (1)}{4}))=i"

"k=2, \sqrt[4]{1}(\cos(\dfrac{0+2\pi (2)}{4})+i\sin(\dfrac{0+2\pi (2)}{4}))=-1"

"k=3, \sqrt[4]{1}(\cos(\dfrac{0+2\pi (3)}{4})+i\sin(\dfrac{0+2\pi (3)}{4}))=-i"

"\sqrt[4]{1}=1"


"\sqrt[4]{1}=i"


"\sqrt[4]{1}=-1"


"\sqrt[4]{1}=-i"